OpenStudy (anonymous):

What is the area of a circle with a radius of 4 meters? Use twenty-two sevenths for pi. twelve and four-sevenths square meters twenty-five and one-seventh square meters fifty and two-sevenths square meters seventy-five and three-sevenths square meters

OpenStudy (anonymous):

The area of a circle is A=pi*(r)^2

OpenStudy (anonymous):

Area=pir^2

OpenStudy (anonymous):

If r=4, do you remember how to find 4^2? I showed you before

OpenStudy (anonymous):

ya we sqaure it

OpenStudy (anonymous):

so what is 4 squared?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

16 time 22/7=?

OpenStudy (anonymous):

Correct! so so far we have A=pi*(16) but they said pi=22/7 so what is 22/7*16?

OpenStudy (anonymous):

ail that should have given you enough. do you think u know what the final answer is?

OpenStudy (anonymous):

on my calculator it says 50.285714

OpenStudy (anonymous):

and that looks (close?) like c

OpenStudy (anonymous):

ur right. so the final answer is..?

OpenStudy (anonymous):

u got it good job:D

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

^What he/she said :P

OpenStudy (anonymous):

lol